On January 1, 2021, Byner Company purchased a used tractor. Byner paid $9,000 down and signed a noninterest-bearing note requiring $35,000 to be paid on December 31, 2023. The fair value of the tractor is not determinable. An interest rate of 12% properly reflects the time value of money for this type of loan agreement. The company’s fiscal year-end is December 31. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Prepare the journal entry to record the acquisition of the tractor.
2. How much interest expense will the company include in its 2021 and 2022 income statements for this note?
3. What is the amount of the liability the company will report in its 2021 and 2022 balance sheets for this note?
Answer #1

Solution 1:

Fair value of tractor = Down payment + Present value of note

= $9,000 + $35,000 * PV factor at 12% for 3rd period

= $9,000 + $35,000 * 0.71178 = $33,912

Journal Entries
Date Particulars Debit Credit
1-Jan-21 Equipment Dr $33,912.00
Discount on notes payable Dr $10,088.00
            To Cash $9,000.00
            To Notes Payable $35,000.00
(To record purchase of tractor)

Solution 2:

Interest expense for 2021 = $24,912 * 12% = $2,989

Interest expense for 2022 = ($24,912 + $2,989) * 12% = $3,348

Solution 3:

Amount of liability the company will report in its 2021 balance sheet for this note = $24,912 + $2,989 = $27,901

Amount of liability the company will report in its 2022 balance sheet for this note = $27,901 + $3,348 = $31,249